Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, pinto, ruggi - 2:13 Saturday 23 October 2021 (53617) Print this report
ISC/TCS: DAS/CH tuning using CITF

The goal of this shift was to perform some tunings of the DAS powers with CITF locked.

All DASes have been switched OFF at 14.38 UTC and the CH powers have been settled to: 

WI CH= 20 mW 

(113 mW on pickoff) 

NI CH =40 mW  

(226 mW on pickoff)

to compensate the cold defects to lock the CITF. At 15.44 UTC the CITF locked.

At 15.56 UTC we started to turn on the WI DAS:
- IN ON at 223 mW. The sidebands started to decrease but no shape change has been observed on B1p. To recover the sideband value, we increased the CH power by 10 mW, from 20 mW→ 30 mW, at 16.22 UTC. As expected the sideband increased. 
- At 16.50 UTC the WI OUT has been switched ON at 600 mW. We observed a worsening of the fringe and we decreased the power from 600 mW → 400 mW at 17.07 UTC. Paolo changed the BS position at 17.38 UTC in TX and TY decreasing the dark fringe power and increasimg the sideband level. In order to understand if the outer ring was the culprit of the BS misalignment, the BS has been moved to the previous working condition and the WI OUT has been switched off at 18.04 UTC. Nothing changed by switching OFF  the outer ring and Paolo optimized again the BS angular positions. At 19.01 UTC, we increased the WI CH power by 10 mW, 30 mW→ 40 mW. The sideband level increased and the dark fringe power decreased.
- At 19.33 UTC the WI OUT has been switched on again (~400 mW) and the tuning of the WI DAS and CH has been stopped.

At 19.58 UTC we started to turn on the NI DAS:
- IN ON at 150 mW. The  sideband power decreased, as observed yesterday. To recover the sideband level, at 20.33 UTC, we increased the NI CH by 20 mW, from 40 mW to 60 mW. The sideband started to increase. At 21.20 UTC, we increased more the NI CH (60 mW → 70 mW).
- At 21.53 UTC, we switched on the NI OUT ON (300 mW) to improve the shape of the fringe. The fringe shape slightly imporved but the sideband level started to decrease. At 22:10, twe increased the power of the NI OUT up to 600 mW in order to further improve the quality of the dark fringe shape. At 22.34 UTC, we slightly decreased the power of the NI IN down to 130 mW. At 22.46 UTC, we increased the power of the NI CH by 10 mW, from 70 mW to 80 mW.

All the main signal during the tuning are shown in attached figure.

The final powers are:

WI CH= 40 mW 

NI CH =80 mW  

WI Inner= 223 mW 

NI Inner =130 mW  

WI Outer=400 mW 

NI Outer = 600 mW 

All the CO2 actuators are injected in the ITF to allow the next lock at CARM null.

We left the ITF with the cavities locked on the IR and the Fmoder tuning disabled from the ITF_lock configuration file.

Images attached to this report
Search Help